inverse of f(x)= cubic root of x+4
Answer by algebrapro18(249)

Notes:
1)The cube root of a number is the same as taking the number to the 1/3 power.
2) f(x) and y are the same thing.
3) Cube root and cube are inverse functions of eachother.

To find the inverse of a function you swap the x and y and solve for y.
So doing so we get:

y = (x+4)^1/3 by my notes above
x = (y+4)^1/3 Cube both sides
x^3 = y+4 Subtract 4 from both sides
x^3-4 = y

So the inverse function of f(x) is f^(-1)(x) = x^3-4.